If one angle of a triangle is 30 degrees more than twice another, and the third angle is equal to the sum of the first two angles, find the measures of each angle. (Hint: the three angles in a triangle add up to 180 degrees)
Answer by ikleyn(52834)

.

x = angular measure of the "another" angle

2x+30 is that of the "one" angle

x + (2x+30) = 3x+30 is that of the third angle.


The basic equation for the sum of the three angles is

    x + (2x+30) + (3x+30) = 180.


Simplify and find x

    6x = 180 - 30 - 30 = 120 degrees.

     x                 = 120/6 = 20 degrees.


ANSWER.  The angles are 20 degs, 2*20+30 = 70 degs and 3*20+30 = 90 degs.


CHECK.  20 + 70 + 90 = 180 degs.   ! correct !

Solved.